### 30 Mar 2026 v1

# Dirac branch-cut modes

Abstract
Bound states arising in Dirac fields are usually attributed to two kinds of features: domain walls where a real Dirac mass field changes sign, which host Jackiw–Rebbi states, and phase singularities in a complex Dirac mass field, which host Jackiw–Rossi zero modes. We show that phase discontinuities, such as branch-cuts of complex branch functions, supply a third distinct binding mechanism. We derive the existence of guided modes that propagate along the cut, called Dirac branch-cut (DBC) modes, which obey an effective one-dimensional relativistic Dirac equation with a reduced mass determined by the phase difference across the cut. When the mass field has fixed magnitude, the DBC modes’ transverse confinement lengths are energy-independent, unlike Jackiw–Rebbi and Jackiw–Rossi states or conventional boundary modes. Using acoustic metamaterials, we realize DBC modes experimentally, and verify their relativistic dispersion, robust transverse confinement length, and ability to propagate along freeform paths. These results show that phase discontinuities in a complex Dirac mass field constitute a versatile design principle for guided modes, with interesting application possibilities for photonic and acoustic metamaterials.
The emergence of bound states in Dirac media is of profound importance for many fields of physics, including particle physics, condensed matter physics, and photonics. For example, the Jackiw–Rebbi state [1], which appears at a domain wall in one-dimensional (1D) space where a Dirac mass field changes sign, is recognized as a theoretical prototype for the edge states of topological insulators in 1D or higher dimensions [2]. Such edge states have many interesting properties, and have been observed in a variety of real materials [3, 4], as well as metamaterials based on photonics, acoustics, and other classical-wave platforms [5, 6, 7, 8, 9, 10]. The concept has also been extended beyond domain walls to features such as higher-order boundaries [11, 12, 13, 14] and spatial defects [15, 16, 17, 18, 19, 20]. Notably, a leading conceptual framework for understanding defect-bound topological states is the Jackiw–Rossi model [21], a generalization of the Jackiw–Rebbi model that features a Dirac field in two-dimensional (2D) space with a complex mass field (composed of two independent Dirac mass parameters); a vortex in the phase of the complex mass binds a localized zero-energy state that is robust against mass field variations that do not alter the winding number [22]. These Jackiw–Rossi states form the conceptual foundation for Majorana excitations in topological superconductors [23, 24], and have recently also been experimentally realized in Kekulé-modulated lattices [25, 26, 27]. In photonics, particularly, there is much ongoing research into using “Dirac vortex” modes for enhancing light-matter interactions, designing new laser cavity designs, and other practical applications [28, 29, 30, 31, 32, 33].
While the effects of domain walls and vortices in Dirac mass fields have been extensively studied, an important intermediate case—a phase discontinuity along a curve—has received little attention. Such discontinuities can occur when the mass is set to a branch of a multi-valued complex function, like a root or logarithm, which contains a branch-cut [34, 35]: i.e., a 1D curve in 2D space along which there is a dislocation, or phase discontinuity, in the complex function. Here, we identify and characterize a new class of bound states arising on such cuts, which we call Dirac branch-cut (DBC) modes. We show that these modes possess a tunable relativistic dispersion, with a reduced effective mass determined by the phase discontinuity. Remarkably, when the magnitude of the mass field (and hence the bulk band gap) are fixed, the DBC modes’ transverse confinement lengths are independent of energy, unlike conventional boundary modes, whose confinement weakens near the band gap edge. These features persist for different cut choices. If the mass field has nonzero winding at infinity, one DBC mode also corresponds to the mid-gap Jackiw–Rossi state [21].
We then realize DBC modes experimentally using classical acoustic structures. Starting from a 2D triangular acoustic crystal, applying Kekulé modulations to the pillar radii [28] maps the acoustic mode amplitudes to the Jackiw–Rossi model. This design provides a great deal of flexibility for setting $m(\mathbf{r})$, allowing for curved and/or bounded cuts. For straight cuts, we experimentally validate the relativistic dispersion relations of the DBC modes, including the variation of the effective mass with the phase discontinuity. We also study DBC modes in other configurations, including line cavities formed by finite cuts ending at fixed branch points [25, 26, 27, 28, 29, 30, 31]. We also demonstrate the potential of DBC modes for freeform waveguiding by showing efficient transmission along a spiral cut with arc length of $\sim 100$ lattice periods.
These results imply that mass phase discontinuities are a useful, and hitherto overlooked, mechanism for guiding Dirac waves. This phase-guiding is distinct from previous schemes based on varying a real Dirac mass [36, 37, 38, 39]. Moreover, while there exist other ways to generate robust freeform waveguides, e.g., using amorphous lattices [40, 41, 43, 42], DBC modes are considerably more straightforward to realize: we simply choose a branch-cut and modulate the lattice accordingly [25, 26, 27, 28, 29, 30, 31], without additional computational steps such as close-packing lattice sites [41, 42]. This ease of realization, together with the unique features of the DBC modes, point to intriguing possibilities for realizing robust waveguides and resonators in acoustics, photonics, and other settings.
We begin by deriving the existence of DBC modes in the Jackiw–Rossi model for a four-component Dirac equation in 2D space [21]. The Hamiltonian is [28]
$H=\tau_{3}\left(-i\sigma_{1}\frac{\partial}{\partial x}-i\sigma_{2}\frac{\partial}{\partial y}\right)+m_{1}(\mathbf{r})\tau_{1}+m_{2}(\mathbf{r})\tau_{2},$ (1)
where the Dirac velocity is normalized to 1, the 2D position vector is $\mathbf{r}\equiv(x,y)$, and the internal Hilbert space is factorized as $\mathbb{C}^{2}\otimes\mathbb{C}^{2}$, with $\sigma_{1,2,3}$ and $\tau_{1,2,3}$ denoting Pauli matrices acting on each factor space. The Clifford algebra allows for two Dirac masses $m_{1},\,m_{2}$, which can be consolidated into a complex mass field $m(\mathbf{r})\equiv m_{1}+im_{2}=|m|\,e^{i\theta}$. Note that $H$ is Hermitian in spite of this complex mass. Jackiw and Rossi considered a specific $m(\mathbf{r})$ with $\theta$ winding by $2\pi$ around the origin, and showed that it hosts a mid-gap bound state [21]; this result was later generalized to an index theorem stating that any $m(\mathbf{r})$ with a $2\pi$ phase winding implies the existence of such a bound state [22].
Figure 1: Properties of Dirac branch-cut (DBC) modes. a, Upper panel: complex plane for the Jackiw–Rossi mass parameter $m$, indicating two points with the same $|m|$ and phase difference $\Delta\theta$. Lower panel: a schematic of a piecewise-constant $m(\mathbf{r})$ distribution in 2D space, constructed from those two $m$ values with phase dislocation along $y=0$. b, Phase dislocation mode (DBC mode) energies versus wavenumber $k_{x}$ for $\Delta\theta=\pi$ (left panel) and $\Delta\theta=2\pi/3$ (middle panel), and versus $\Delta\theta$ at $k_{x}=0$ (right panel). Cyan dashes show the results obtained analytically using Eqs. (4)–(5). Color maps show the minimum eigenvalue magnitude of the scattering matrix, computed directly from the Hamiltonian (1); a zero value corresponds to a boundary state (see Supplementary Information). The bulk bands are shaded in gray. c, Phase profile of a mass distribution made from a branch function $m=(z-3)^{1/2}(z+3)^{1/2}$, where $z=x+iy$. d, Numerically-obtained DBC mode energies versus mode order for the mass distribution in c. e, Spatial profiles for two of the modes in d: the Jackiw–Rossi state at $E=0$ (left panel), and the order-4 mode (right panel). f, Phase profile of a “phase-only” mass distribution $m=m_{0}\mathrm{arg}[(z-3)^{1/2}(z+3)^{-1/2}]$, with $m_{0}=5$. g, DBC mode energies for the mass distribution in f (blue markers), and for another mass distribution with $\Delta\theta=2\pi/3$ using the same cut (yellow markers). The theoretically-predicted minigap for the latter is indicated by horizontal dashes. h Representative spatial profiles of two of the DBC modes for $\Delta\theta=\pi$. All the DBC modes have very similar confinement lengths, up to high orders approaching the bulk bands.
Now, instead of phase singularities, we study phase dislocations. We first consider the piecewise-constant mass field in Fig. 1a, with $m=m_{0}e^{i\theta_{+}}$ for $y>0$ and $m=m_{0}e^{i\theta_{-}}$ for $y<0$, where $m_{0}\in\mathbb{R}^{+}$ is a constant. In each domain, there is a bulk gap $m_{0},m_{0}]$, while at the domain wall $y=0$ there is a phase discontinuity $\Delta\theta=\theta_{+}-\theta_{-}$. For any $\Delta\theta\neq 0$, there exist DBC modes along the domain wall, with the Jackiw–Rebbi-like form
$\psi(y)=\begin{cases}e^{+\kappa y}\,e^{-i\tau_{3}\theta_{-}/2}\,|\chi_{-}\rangle,&y<0,\\ e^{-\kappa y}\,e^{-i\tau_{3}\theta_{+}/2}\,|\chi_{+}\rangle,&y>0,\end{cases}$ (2)
where $\kappa=\left(m_{0}^{2}+k_{x}^{2}-E^{2}\right)^{1/2}$ is the skin depth, $k_{x}$ is the wavenumber along $x$, and the energy $E$ lies in the bulk gap ($|E|<m_{0}$), and $|\chi_{\pm}\rangle$ are spinors satisfying
$\big[\tau_{3}\left(k_{x}\sigma_{1}\pm i\kappa\sigma_{2}\right)+m_{0}\tau_{1}\big]|\chi_{\pm}\rangle=E|\chi_{\pm}\rangle.$ (3)
It can be shown that the continuity of $\psi(y)$ along $y=0$ constrains the solutions to Eq. (3) to a $\Delta\theta$-dependent subspace (see Methods). The DBC mode eigenproblem then maps to a two-component 1D Dirac equation, with
$E^{2}$ $=k_{x}^{2}+m_{b}^{2},\;\;\;m_{b}=m_{0}\left|\cos(\Delta\theta/2)\right|,$ (4)
$\kappa$ $=m_{0}\left|\,\sin(\Delta\theta/2)\,\right|.$ (5)
In Eq. (4), the effective mass $m_{b}$ depends on $m_{0}$ and the phase dislocation $\Delta\theta$. For $\Delta\theta=\pi$, which corresponds to a Jackiw–Rebbi-like sign flip of $m$ across the domain wall, we have $m_{b}=0$; for other $\Delta\theta$, we have $0<m_{b}<m_{0}$ (i.e., there is a minigap enclosed within the bulk gap). The addition of a phase gradient $v_{\theta}y$ retains this result, but with a correction to $m_{b}$ (see Methods). Importantly, because $\kappa$ is independent of $k_{x}$, all the DBC modes have the same transverse confinement length, regardless of where they lie in the bulk gap. This property is due to the 1D dispersion relation (4) and the bulk dispersion relation for guided modes, $E^{2}=k_{x}^{2}-\kappa^{2}+m_{0}^{2}$, from Eq. (1); these involve the same Dirac velocity, and combining them leads to Eq. (5).
In Fig. 1b, we compare the analytic result (4) to numerical results obtained directly from Eq. (1), obtaining exact agreement for the dispersion relations (left and middle panels) and their $\Delta\theta$-dependence (right panel).
From these findings, we hypothesize that the DBC modes arise chiefly from the presence of a discontinuity in $\theta(\mathbf{r})$, and will also exist for phase dislocations along arbitrary 1D curves. A convenient way to generate phase dislocations is to use a branch-cut—a 1D curve connecting different Riemann sheets of a complex multi-valued function, like a complex root or logarithm [34]. A branch function is specified by choosing one Riemann sheet, and is differentiable for all $\mathbf{r}$ except at the cut, where there is a constant phase discontinuity given by the monodromy (i.e., the multiplicity of the Riemann sheets). As an example, take $f(z)=(z-z_{1})^{\alpha}(z-z_{2})^{\beta}$, where $z=x+iy$, and $\alpha,\beta\in\mathbb{Q}$ with $\alpha+\beta\in\mathbb{Z}$. The branch points are $z_{1}$ and $z_{2}$, with no branch point at infinity; we can place the cut arbitrarily, so long as it does not cross itself and terminates at the branch points. All along the cut, there is a phase discontinuity of $\Delta\theta\equiv 2\pi\alpha\equiv-2\pi\beta\,(\textrm{mod}~2\pi)$.
In Fig. 1c, we show the distribution of $\theta(\mathbf{r})=\mathrm{arg}[m(\mathbf{r})]$ for $m(\mathbf{r})=(z-3)^{1/2}(z+3)^{1/2}$. The cut is chosen to be a sinusoidal curve connecting the branch points at $z=\pm 3$, and the phase discontinuity is $\Delta\theta=\pi$. Since $|m|\rightarrow\infty$ as $|\mathbf{r}|\rightarrow\infty$, there are no free solutions, only bound states. We solve for these numerically using the finite element method (FEM) (see Methods), obtaining the spectrum in Fig. 1d, with discrete energies extending to $E\rightarrow\pm\infty$. As the branch function has a phase winding of $2\pi$ upon encircling both branch points, the index theorem for the Jackiw–Rossi model predicts a mode at $E=0$ [21, 22], consistent with our obtained spectrum; in other words, the Jackiw–Rossi state is one of the DBC modes. However, all the DBC modes are strongly bound to the cut, with similar confinement lengths; two of them, the zero mode and the order-4 mode, are plotted in Fig. 1e.
Figure 2: Implementation of DBC modes in an acoustic crystal. a, Schematic of the acoustic structure, consisting of solid pillars (pink circles) surrounded by air. The pillar positions are fixed, and their radii are modulated according to Eq. (6); in this plot, we take $\theta=-0.35\pi$. Dashed and solid circles indicate the radii before and after modulation, respectively. b, A phase map $\theta(\mathbf{r})$ with a $\pi$ phase dislocation along $y=0$ (left panel), and the corresponding acoustic structure (right panel). c, DBC mode dispersion for $\Delta\theta=\pi$ (left panel) and $\Delta\theta=2\pi/3$ (right panel). Color maps show the Fourier transformed excitation spectrum obtained from experiments, while cyan dashes show the dispersion relations from FEM simulations of the acoustic structure (see Methods). The bulk bands are shaded in gray. Markers indicate four frequencies between 10.7 kHz–11.3 kHz (with 0.2 kHz spacing) used in subsequent subplots. d, Experimental setup for measuring spatially-resolved excitation spectra. The location of the acoustic source is indicated by the loudspeaker symbol. e, Measured profiles for DBC modes at 10.9 kHz (upper panel) and 11.3 kHz (lower panel). f, Semilogarithmic plot of the transverse acoustic pressure profiles, measured along a line of constant $x$ with the four frequencies marked in c. These experimental results are normalized to the value at $y=0$, and agree with the frequency-independent theoretical prediction derived from the Jackiw–Rossi model (solid green line).
We can also devise mass distributions with finite bulk gaps, which are more suited to experimental realization (as discussed below). An expedient way to do this is to take only the phase of a complex branch function, keeping the magnitude constant. As an example, we take $|m|=5$ together with $\theta=\mathrm{arg}\{[(z-3)/(z+3)]^{1/2}\}$, using an arc-like cut, as shown in Fig. 1f; note that the non-holomorphic nature of $m(\mathbf{r})$ does not invalidate the preceding discussion. There is now a bulk gap $E\in[-5,5]$, with free states outside and DBC modes inside the gap (Fig. 1g, blue circles). The net phase winding is zero in this case, so the DBC modes do not include a Jackiw–Rossi mode. For comparison, we also plot the spectrum for $|m|=5$ and $\theta=\mathrm{arg}\{[(z-3)/(z+3)]^{1/3}\}$, which has phase discontinuity $\Delta\theta=2\pi/3$ (Fig. 1g, yellow circles). These DBC modes exhibit a minigap closely matching the one derived from Eq. (4), as indicated by dashes in the figure. The DBC modes are again strongly confined to the cut, as shown in Fig. 1h. Remarkably, even high-order DBC modes very close to the bulk band edge (e.g., mode # 10 in Fig. 1g,h) have the same tight confinement, consistent with the prediction from Eq. (5).
In the Supplementary Information, we show the mode profiles for other DBC modes, whose characteristics are found to be very close to those in Fig. 1e,h. We also study other mass distributions, such as one where the DBC minigap is nonzero while the overall phase winding is nonzero; in that case, there exists a Jackiw–Rossi state that is distinct from the DBC modes, lying within the minigap and localized near one of the branch points.
The Jackiw–Rossi model has previously been realized in several classical-wave experiments [25, 26, 27, 28, 29, 30, 31], and although those works focused on phase singularities and the associated Jackiw–Rossi modes (also called “Dirac vortex states” in this context), we can leverage on such ideas to realize DBC modes. We choose an acoustic crystal design based on the work of Torrent and Sánchez-Dehesa [44]. Starting from a triangular lattice of pillars of equal radius $r_{0}$ and next-nearest-neighbor distance $a$, we apply the modulation shown in Fig. 2a: each pillar is kept at a fixed center position $\mathbf{r}$, but its radius is set to
$R(\textbf{r})=R_{0}+\Delta R\,\mathrm{cos}\left[\textbf{K}\cdot\textbf{r}+\theta(\textbf{r})\right],$ (6)
where $\textbf{K}=\textbf{K}_{+}-\textbf{K}_{-}$, with $\textbf{K}_{\pm}=\left[\pm 4\pi/\sqrt{3}a,0\right]$ denoting the wave vectors at the corners of the original lattice’s Brillouin zone [26]. The modulation couples the Dirac cones, and in the long wavelength limit the four mode amplitudes (two Dirac cones, with two amplitudes each) are governed by the Jackiw–Rossi Hamiltonian (1), with $\Delta R\leftrightarrow|m|$ [28]. In our experiments, we take $a=29$ mm and $R_{0}=0.2a$, which yields a Dirac frequency of 10.4 kHz for the unmodulated lattice (see Supplementary Information, Fig. S3). We also set $\Delta R=0.13d$, which corresponds to keeping $|m|$ constant and varying only the phase.
Figure 3: Coupled waveguide-cavity system based on DBC modes. a, Upper panel: diagram $\theta(\textbf{r})$ created from branch function $f(z)=(z-150)^{1/2}(z+150)^{-1/2}$, with $z=x+i4y$. Lower panel: the acoustic crystal modulated with the diagram $\theta(\textbf{r})$. Blue dashed lines indicate the positions of cuts I and $I$, which are separated by a distance of 7070 mm. Scale bar: 100 mm. b, Upper panel: acoustic power transmittance to the right side of cut II. Middle panel: transmittance spectrum from FEM simulations. Lower panel: quality (Q) factors of the DBC modes on cut I. The bulk bands are shaded in gray. c, The profiles of DBC modes from experiment (upper panel) and simulations (lower panel) at the frequency marked by the star in b.
We begin by implementing a sample with an approximately piecewise-constant phase distribution with $\Delta\theta=\pi$, analogous to Fig. 1a. Fig. 2b shows the phase distribution (left panel) and the modulated acoustic lattice (right panel). We experimentally determine the excitation spectrum by applying an acoustic source and measuring the acoustic pressure at different points along the cut (Fig. 2c,d), followed by a Fourier transform (see Methods). In Fig. 2c, the color maps show the resulting excitation spectrum for this acoustic sample, as well as another similar sample with $\Delta\theta=2\pi/3$. Both sets of data agree quantitatively with the dispersion curves calculated directly from the acoustic equations via FEM (see Methods), plotted as cyan dashes. These dispersion curves are approximately hyperbolic near the center of the bulk gap, in accordance with the theory, but deviate from the hyperbolic form as they approach the bulk bands. For $\Delta\theta=\pi$, the DBC minigap vanishes and agrees with the prediction in Fig. 1b. For $\Delta\theta=2\pi/3$, the minigap approximately matches the theoretical value of $2m_{b}$ calculated from Eq. (4) using the observed bulk gap.
We now measure the spatial profiles of the DBC modes in the frequency regime where the dispersion relation is hyperbolic. The results are plotted in Fig. 2e,f, and reveal exponential decays with distance $y$ from the cut. In Fig. 2f, we include the transverse profile predicted by the theoretical model (with $m_{0}$ again fitted using the observed bulk gap, and the Dirac velocity fitted to the slope of the dispersion curve in Fig. 2c), which has decay length $\kappa^{-1}\approx 30$ mm. Over the measured frequency range, which spans a large portion of the bulk gap, the experimental profiles agree well with the theoretical line.
Next, to illustrate the flexibility of the DBC modes, we study samples with different cut configurations.
In Fig. 3a, the upper panel shows a phase distribution constructed from a bounded cut (I) adjacent to an unbounded cut (II), forming a waveguide-cavity system like those found in integrated photonics [41, 42, 46]. The DBC modes on cut I turn into a set of discrete modes for a line cavity (i.e., a Fabry-Pérot-like resonator), while the DBC modes on cut II are waveguide modes that couple to the cavity. A photograph of the experimental sample is shown in the lower panel of Fig. 3a. We excite cut II from the left, and monitor the acoustic power transmitted to its other end. Within the bulk gap, the measured transmittance spectrum (Fig. 3b, upper panel) exhibits a dip near mid-gap that may be attributed to local symmetry-breaking along the domain wall [45] (as well as the reduction in group velocities above and below the minigap). We also observe a sequence of narrow dips, which are absorption resonances caused by critical coupling to the cavity [46]. These results are in agreement with the more ideal transmittance spectrum produced by FEM simulations (Fig. 3b, middle panel), including quantitative matches for the resonance frequencies (vertical dashes). Furthermore, we numerically calculate the eigenmodes for the line cavity in the absence of cut II, and find that the eigenfrequencies closely match the absorption dips (Fig. 3b, bottom panel). These simulation results also show that the eigenmodes’ quality (Q) factors (based on radiative losses at the edges of the computational cell) remain approximately constant, or increase slightly, as the eigenfrequencies approach the upper bulk band. In Fig. 3c, we present experimental measurements (upper panel) and simulations (lower panel) for the field intensity profile at one of the absorption resonances. The excited line cavity mode is strongly confined to cut I, as expected. In the Supplementary Information, we perform a comparison to an analogous line cavity based on conventional defect modes, and find that the DBC-based cavity modes indeed exhibit less frequency variation.
Figure 4: DBC modes on a spiral. a, A modulation phase profile $\theta(\textbf{r})$ with a spiral-shaped cut. b, Photograph of the acoustic lattice generated from this phase profile, with the cut indicated by the cyan curve. The location of the acoustic source in the experiment is indicated by the loudspeaker symbol. c, The acoustic pressure distribution measured at 10.95 kHz. d, Spatial profiles measured along a line transverse to the spiral (dashes in c) at two different frequencies (blue dots and red circles). The solid green line is the theoretical prediction.
A major advantage of DBC modes is their inherent flexibility: they can transmit along arbitrary paths that can be easily generated by assigning branch-cuts. To demonstrate this, we design a spiral-shaped cut with $\Delta\theta=\pi$, as shown in Fig. 4a. Keeping $|m|$ constant, we use the same scheme as before to generate the acoustic structure. Fig. 4b shows simulation results for the transmittance across a segment of the cut (between the yellow dashes Fig. 4a), in a frequency range corresponding to the upper branch of DBC modes (see Fig. 2d). The fabricated sample is shown in Fig. 4c (left panel). Placing an acoustic source at the center of the spiral, we probe the acoustic pressure distribution at regularly-spaced points across the sample. Evidently, the acoustic waves emitted by the source propagate outward along the spiral, with no sign of Anderson localization in the longitudinal direction. Fig. 4d shows the field profiles along a line perpendicular to the spiral; the results measured at two representative frequencies exhibit confinement lengths consistent with the continuum theory (solid green line). Moreover, we note that the DBC modes appear to be at least as performant as the freeform topological waveguides generated using amorphous lattices [41, 42], but are easier to generate as there is no need to generate a lattice by computational close-packing.
In summary, we have proposed and implemented a new class of robust guided modes in Dirac media called Dirac branch-cut (DBC) modes, which occur along 1D phase discontinuities and are distinct from both Jackiw–Rebbi [1] and Jackiw–Rossi [21] bound states. We have proven theoretically, and verified experimentally, that DBC modes obey a 1D relativistic dispersion relation. We have presented an explicit acoustic realization; this design is easily transferrable to photonics (see Supplementary Information). We have verified that the acoustic DBC modes can form line cavities and propagate robustly along complicated paths. In future, we expect that the lattice modulation scheme can be further refined to bring the actual DBC modes closer to their ideal counterparts, including making the dispersion relations hyperbolic over a larger portion of the bulk gap. The simple realization scheme makes DBC modes versatile candidates for photonic (see Supplementary Information) or acoustic waveguides, and their strong confinement could be helpful for managing inter-waveguide crosstalk. Thus, our results point to a promising pathway for practical device applications based on the fundamental properties of relativistic band structures.
Methods
Solution for a straight phase dislocation
Starting from the Hamiltonian (1), we let the mass distribution be invariant along $x$, i.e., $m_{1}+im_{2}=m(y)$. We aim to derive the boundary states for a phase dislocation along the line $y=0$. Introducing the wavenumber $k_{x}$, the Schrödinger equation reduces to
$\left[\tau_{3}\left(\sigma_{1}k_{x}-i\sigma_{2}\partial_{y}\right)+m_{0}(y)\,U_{\theta}^{\dagger}\tau_{1}U_{\theta}\right]\psi=E\,\psi,$ (7)
where $U_{\theta}\equiv e^{i\theta\tau_{3}/2}$, and $\psi(y)$ is a four-component wavefunction. We now define
$\psi(y)=U^{\dagger}_{\theta(y)}\,\varphi(y).$ (8)
When this is plugged into (7), the $\partial_{y}$ gives rise to an additional gauge term because $\theta$ depends implicitly on $y$:
$\partial_{y}\left[U^{\dagger}_{\theta}\varphi(y)\right]$ $=U^{\dagger}_{\theta}\left[\partial_{y}-\frac{i\theta^{\prime}}{2}\,\tau_{3}\right]\,\varphi(y),$ (9)
where $\theta^{\prime}\equiv d\theta/dy$. Hence, Eq. (7) simplifies to
$\left[\tau_{3}\sigma_{1}k_{x}-i\tau_{3}\sigma_{2}\partial_{y}-\frac{1}{2}\sigma_{2}\,\theta^{\prime}+m_{0}(y)\tau_{1}\right]\varphi=E\varphi.$ (10)
Let us now choose the mass function
$m(y)=m_{0}\,e^{iv_{\theta}y}\,\times\begin{cases}e^{i\theta_{+}},&y>0\\ e^{i\theta_{-}},&y<0,\end{cases}$ (11)
where $m_{0}\in\mathbb{R}^{+}$ and $v_{\theta}\in\mathbb{R}$. This has a phase discontinuity of $\Delta\theta=\theta_{+}-\theta_{-}$ at $y=0$. The parameter $v_{\theta}$ gives a uniform phase gradient, with $v_{\theta}=0$ corresponding to the piecewise-constant mass function in Fig. 1a. Hence,
$\theta^{\prime}=v_{\theta}+\Delta\theta\,\delta(y).$ (12)
We now take the ansatz
$\varphi(y)=\begin{cases}|\chi_{+}\rangle\,e^{-\kappa y},&y>0\\ |\chi_{-}\rangle\,e^{\kappa y},&y<0,\end{cases}$ (13)
where $\kappa\in\mathbb{R}^{+}$ and $|\chi_{\pm}\rangle$ are four-component spinors. Plugging this into (10), for $y\neq 0$, gives
$A_{\pm}|\chi_{\pm}\rangle$ $=E|\chi_{\pm}\rangle,$ (14)
$A_{\pm}$ $\equiv k_{x}\tau_{3}\sigma_{1}\pm i\tau_{3}\sigma_{2}\kappa-\frac{v_{\theta}}{2}\,\sigma_{2}+m_{0}\tau_{1}.$ (15)
To handle the boundary conditions at $y=0$, we require the original wavefunction $\psi(y)$ to be continuous at $y=0$, so the rotated wavefunction $\varphi(y)$ is discontinuous. Integrating (10) across an infinitesimal interval over $y=0$ using Eq. (12), we obtain
$-i\tau_{3}\sigma_{2}\left[\varphi(0^{+})-\varphi(0^{-})\right]$ $=\frac{\Delta\theta}{4}\sigma_{2}\left[\varphi(0^{+})+\varphi(0^{-})\right].$
Referring to (13), this is satisfied by the boundary condition $|\chi_{+}\rangle=U_{\Delta\theta}\;|\chi_{-}\rangle$. We can plug this into Eq. (14) to eliminate (say) $\chi_{+}$, obtaining
$\left(U_{\Delta\theta}^{\dagger}A_{+}U_{\Delta\theta}-A_{-}\right)|\chi_{-}\rangle=0.$ (16)
Using (15), the terms in parentheses can be simplified to
$2i\kappa\tau_{3}\sigma_{2}+m_{0}\left[\tau_{1}(\cos\Delta\theta-1)+\tau_{2}\sin\Delta\theta\right].$
What is the spectrum of this operator? Pick an eigenstate of $\sigma_{2}$, i.e., $|\chi_{-}\rangle=|u\rangle\otimes|v\rangle$ where $\sigma_{2}|v\rangle=\pm|v\rangle$. Then (16) reduces to
$\begin{pmatrix}\pm 2i\kappa&m_{0}\left(e^{-i\Delta\theta}-1\right)\\ m_{0}\left(e^{i\Delta\theta}-1\right)&\mp 2i\kappa\end{pmatrix}|u\rangle$ $=0.$ (17)
For either choice of $\pm$, the determinant vanishes when
$4\kappa^{2}-m_{0}^{2}\;\Big|\cos\Delta\theta-1-i\sin\Delta\theta\Big|^{2}=0.$ (18)
Hence, we arrive at Eq. (5).
Note that the result for $\kappa$ solves (17) for both choices of $\pm$, i.e., both eigenstates of $\sigma_{2}$. It follows that $|\chi_{-}\rangle$ is a superposition of these two choices. To be specific, let us take the eigenstates of $\sigma_{2}$ to be $|v_{+}\rangle=2^{-1/2}[1,i]^{T}$ and $|v_{-}\rangle=\sigma_{1}|v_{+}\rangle$. In each case, the non-Hermitian matrix in (17) has equal and opposite eigenvalues that are either purely real or imaginary. The zero eigenvalue condition is an exceptional point where the matrix has only one eigenvector, which we denote by $|u_{+}\rangle$ or $|u_{-}\rangle$. Evidently, $|u_{+}\rangle$ and $|u_{-}\rangle$ depend only on $\Delta\theta$ (and are independent of $m_{0}$ and $k_{x}$), and can be defined such that
$|u_{+}\rangle=\frac{1}{\sqrt{2}}\begin{pmatrix}\textrm{sgn}[\sin(\Delta\theta/2)]\\ e^{i\Delta\theta/2}\end{pmatrix},\quad|u_{-}\rangle=\tau_{3}|u_{+}\rangle.$ (19)
Note that $\langle u_{+}|u_{-}\rangle=0$. With these, the eigenproblem for $|\chi_{-}\rangle$ is converted into a restricted eigenproblem (i.e., one where the solution is constrained to a subspace):
$\left[k_{x}\tau_{3}\sigma_{1}-i\tau_{3}\sigma_{2}\kappa-\frac{v_{\theta}}{2}\sigma_{2}+m_{0}\tau_{1}\right]$ $|\chi_{-}\rangle=E|\chi_{-}\rangle,$ (20)
where $\kappa$ is given by (5) and
$|\chi_{-}\rangle$ $=\alpha\,|u_{+}\rangle\,|v_{+}\rangle+\beta\,|u_{-}\rangle\,|v_{-}\rangle$ (21)
for some $\alpha,\beta\in\mathbb{C}$. Applying either $\langle u_{+}|\,\langle u_{+}|$ or $\langle u_{-}|\,\langle u_{-}|$ to the left side, we obtain a pair of scalar equations for $\alpha$ and $\beta$ that can be consolidated into
$\Big(m_{b}\,\sigma_{3}+k_{x}\sigma_{1}\Big)\begin{pmatrix}\alpha\\ \beta\end{pmatrix}=E\begin{pmatrix}\alpha\\ \beta\end{pmatrix}$, (22)
$m_{b}=m_{0}\;\textrm{sgn}[\sin\Delta\theta]\,\Big|\cos(\Delta\theta/2)\Big|-\frac{v_{\theta}}{2}.$ (23)
This 1D Dirac equation, with effective Dirac mass $m_{b}$, gives the dispersion relation (4). The spectral theorem now implies that the restricted eigenproblem (20) has two solutions for each $k_{x}$, which is significant as restricted eigenproblems are not guaranteed to possess solutions. This clarifies why the $4\times 4$ Dirac equation exhibits two, not four, boundary states for each $k_{x}$, and retroactively justifies the ansatz (13).
Numerical solutions
The results presented in Fig. 1ch are obtained by numerically solving the Jackiw–Rossi equations using the FEM software package COMSOL Multiphysics with custom-defined partial differential equations expressed in weak form (see Supplementary Information). In these calculations, the cut is located along the curve $y=A\mathrm{sin}[N\pi(x-z_{1})/(z_{2}-z_{1})]$, where $z_{1}=-3$, $z_{2}=3$, and $A=(z_{2}-z_{1})/6$. We take $N=2$ in Figs. 1ce, and $N=1$ in Figs. 1fh.
The numerical simulations of the acoustic crystal, such as those shown in Fig. 2c and Fig. 3bc, are also done using COMSOL Multiphysics. The simulations solve the acoustic equations outside the pillars, using air density 1.18 $\mathrm{kg}\,\mathrm{m}^{-3}$ and speed of sound 343 $\mathrm{ms}^{-1}$; the pillar surfaces are treated as hard boundaries, which is valid due to the large impedance difference between air and the resin material. Due to the limited acoustic pressure in our experiment, nonlinearities are neglected [47].
When calculating the dispersion relations for the DBC modes (Fig. 2c), we use a supercell with periodic and radiation conditions on the boundaries parallel to $y$ and $x$, respectively. When calculating the transmitted power and mode profiles in Figs. 3 and 4, radiation conditions are applied on all edges of the computational cell. We also use the numerical model to calculate the band structure of the acoustic structure, and validate that the Dirac mode amplitudes are governed by the Jackiw–Rossi model as described in Ref. [44]; for details, see the Supplementary Information.
Sample fabrication and characterization
Each acoustic sample is composed of an upper plate, lower plate, and sealing plug. The upper plate is made of transparent acrylic with modulus 3000 MPa and density $1190\,\mathrm{kg}/\mathrm{m}^{3}$, with holes at designated positions for inserting acoustic probes. The lower plate and sealing plug are manufactured via stereolithography, using photosensitive resin with modulus 2880 MPa and density 1100 $\mathrm{kg}/\mathrm{m}^{3}$. The pillars are fabricated directly on the lower plate.
For the dispersion measurements of Fig. 2c, we set up an acoustic source (Brüel &\& Kjær Type 2735) to emit a monochromatic signal through a narrow tube (radius $\approx 1.5\,\textrm{mm}$ and length $100\,\textrm{mm}$) to a point on the sample’s left edge. The signal is picked up by two probe microphones (Brüel &\& Kjær Type 4182). The first microphone scans over 48 small holes (diameter $2\,\textrm{mm}$, step size $12.56\,\textrm{mm}$) placed along the line of the phase dislocation (parallel to the $x$ direction). During each measurement, the other holes are filled with sealing plugs to prevent acoustic leakage. The second microphone is fixed towards the source to pick up a reference signal. The acoustic pressure at different positions is recorded with a data collector (Brüel&\&Kjær Type 3560-C), and a Fourier transform is performed to obtain the $k_{x}$-dependence. This process is repeated at different frequencies to obtain the results in Fig. 2c.
For the intensity profile measurements of Figs. 3 and 4, we use a monochromatic acoustic source and two probe microphones in a similar manner. The acoustic signal is picked up along a grid of holes with spacing $16.74\,\textrm{mm}$ in both the $x$ and $y$ directions.
Acknowledgements
This work was supported by the Singapore National Research Foundation (NRF) under Competitive Research Program (CRP) Nos. NRF-CRP23-2019-0005, NRF-CRP23-2019-0007, and NRF-CRP29-2022-0003, and the NRF Investigatorship NRF-NRFI08-2022-0001; the Singapore Ministry of Education Singapore (MOE-T2EP50122-0019); the Singapore A*STAR Grant No. A2090b0144; the China Scholarship Council: 202406280469.
Supplemental Materials for
“Dirac branch-cut modes”
Bofeng Zhu, Chengzhi Ma, Qiang Wang, Gui-Geng Liu,
Xiuhai Zhang, Qi Jie Wang, Baile Zhang, and Y. D. Chong
S1 Higher-order DBC modes
In Fig. 1h of the main text, we plotted the spatial profiles for several Dirac branch-cut (DBC) modes on an exemplary bounded cut. Here, we present the spatial profiles for several other modes. In Fig. S1a, we plot the profiles for DBC modes # 4 and # 7 at $\Delta\theta=\pi$ (all other parameters are the same as in Fig. 1h of the main text). It can be seen that the mode profiles are extremely similar to those of modes # 1 and # 10 from in the main text. In Fig. S1b, we plot all four modes’ profiles along the center line $x=0$. This reveals that the confinement lengths are all highly similar, consistent with the analytic prediction derived from straight cuts.
Figure S1: DBC modes on a bounded cut. a, Spatial profiles of DBC modes on a curved cut with $\Delta\theta=\pi$. All the other parameters are the same as in Fig. 1h of the main text. b, Profiles along the line $x=0$ for the two DBC modes in a, as well as the two modes plotted in Fig. 1h of the main text. All four modes have very similar confinement lengths. cd, Mode profiles for $\Delta\theta=2\pi/3$.
In Fig. S1c, we plot the spatial profiles of two representative modes for the $\Delta\theta=2\pi/3$ case. The corresponding profiles along $x=0$ are plotted in Fig. S1d. Once again, the modes at very different energies have the same confinement lengths, averaging to $\kappa\approx 0.87$, while the analytic value [Eq. (5) of the main text] is $\kappa=\sin(\pi/3)/\sin(\pi/2)\approx 0.86603$.
In Fig. 1ce of the main text, we presented a case where the Jackiw–Rossi state coincides with one of the DBC modes, and has properties similar to all the other DBC modes. It is also possible to design $m(\mathbf{r})$ functions hosting Jackiw–Rossi modes distinct from DBC modes. For example, Fig. S2a shows the phase distribution for the mass field $m(\mathbf{r})=(z-3)^{1/3}(z+3)^{2/3}$. This has a net winding of $2\pi$ at large distances, and therefore hosts a Jackiw–Rossi state at $E=0$. At the same time, the phase discontinuity at the cut is $2\pi/3$, so the DBC modes are gapped—i.e., their energies cannot overlap with that of the Jackiw–Rossi state, as shown in Fig. S2b. The spatial profiles for several modes are plotted in Fig. S2cd. In this case, we find that the Jackiw–Rossi mode does not spread out along the cut, as the DBC modes do, but is localized near one of the ends.
Figure S2: Separation of Jackiw–Rossi states and DBC modes. a, Phase map for the mass field $m(\mathbf{r})=(z-3)^{1/3}(z+3)^{2/3}$, using a sinusoidal half-period (with amplitude 1) for the cut. b, Eigenenergy spectrum, showing the Jackiw–Rossi state at $E=0$ and a sequence of DBC modes with a finite gap. cd, Spatial profiles for the Jackiw–Rossi mode, and three representative DBC modes indicated by the corresponding markers in b.
S2 Scattering matrix calculations for straight phase dislocations
In this section, we describe a direct numerical solution for the Jackiw-Rossi model for a straight semi-infinite phase dislocation, which is helpful for checking the analytical derivations (see Fig. 1b in the main text). Starting from the $4\times 4$ Dirac equation in 2D, we replace $i\partial/\partial x\rightarrow k_{x}$ and write the Hamiltonian as
$H=\begin{bmatrix}0&k_{x}-\partial_{y}&m(y)&0\\ k_{x}+\partial_{y}&0&0&m(y)\\ m^{*}(y)&0&0&-k_{x}+\partial_{y}\\ 0&m^{*}(y)&-k_{x}-\partial_{y}&0\\ \end{bmatrix}$. (S1)
For simplicity, we assume the piecewise-constant $m$ function shown in Fig. 1a of the main text (i.e., $v_{\theta}=0$ and $\theta_{-}=0$). Then (S1) can be written as
$H$ $=H_{0}(m,k_{x})+\Lambda\frac{\partial}{\partial y},$ (S2)
$H_{0}(m,k_{x})$ $=\begin{bmatrix}0&k_{x}&m&0\\ k_{x}&0&0&m\\ m^{*}&0&0&-k_{x}\\ 0&m^{*}&-k_{x}&0\\ \end{bmatrix},\;\;\;\Lambda=\begin{bmatrix}0&-1&0&0\\ +1&0&0&0\\ 0&0&0&+1\\ 0&0&-1&0\\ \end{bmatrix}=-\Lambda^{-1}.$ (S3)
First, consider $y\leq 0$. Taking $\theta_{-}=0$ for simplicity (i.e., $m=m_{0}$), we seek solutions for fixed $E$ of the form
$\psi(y)=e^{\gamma y}|\mu\rangle,\quad H\psi(y)=E\psi(y).$ (S4)
Plugging this into (S2)–(S3), and doing some rearrangement, yields
$\Lambda\Big[H_{0}(m_{0},k_{x})-E\Big]|\mu\rangle=\gamma|\mu\rangle.$ (S5)
It can be shown that this eigenproblem’s eigenvalues come in pairs $\{\pm\gamma_{1},\pm\gamma_{2}\}$, which can be either real (evanescent waves) or imaginary (propagating waves). We use the ++ label for evanescent waves that grow to the right ($\gamma\in\mathbb{R}^{+}$) as well as propagating waves with positive phase velocity ($\gamma=ik$ where $k\in\mathbb{R}^{+}$); and conversely for the - label. The corresponding eigenvectors are denoted by $\{|\mu_{1}^{\pm}\rangle,|\mu_{2}^{\pm}\rangle$. Hence, the wavefunction has the form
$\psi(y)=a_{1}e^{\gamma_{1}y}|\mu_{1}^{+}\rangle+a_{2}e^{\gamma_{2}y}|\mu_{2}^{+}\rangle+b_{1}e^{-\gamma_{1}y}|\mu_{1}^{-}\rangle+b_{2}e^{-\gamma_{2}y}|\mu_{2}^{-}\rangle,\;\;\;(y\leq 0),$ (S6)
for some coefficients $a_{1(2)},b_{1(2)}\in\mathbb{C}$.
The procedure for $y\geq 0$ is similar, but with $m=m_{0}e^{i\Delta\theta}$. After constructing the eigenproblem analogous to (S5), we obtain a second set of eigenvalues $\{\pm\eta_{1},\pm\eta_{2}\}$, with eigenvectors $\{|\nu_{1}^{\pm}\rangle,|\nu_{2}^{\pm}\rangle\}$. The wavefunction has the form
$\psi(y)=c_{1}e^{\eta_{1}y}|\nu_{1}^{+}\rangle+c_{2}e^{\eta_{2}y}|\nu_{2}^{+}\rangle+d_{1}e^{-\eta_{1}y}|\nu_{1}^{-}\rangle+d_{2}e^{-\eta_{2}y}|\nu_{2}^{-}\rangle,\;\;\;(y\geq 0),$ (S7)
for $c_{1,2},d_{1,2}\in\mathbb{C}$. We now impose continuity of the four-component wavefunction, by equating (S6) to (S7) at $y=0$:
$a_{1}|\mu_{1}^{+}\rangle+a_{2}|\mu_{2}^{+}\rangle+b_{1}|\mu_{1}^{-}\rangle+b_{2}|\mu_{2}^{-}\rangle$ $=c_{1}|\nu_{1}^{+}\rangle+c_{2}|\nu_{2}^{+}\rangle+d_{1}|\nu_{1}^{-}\rangle+d_{2}|\nu_{2}^{-}\rangle$ (S8)
$\Rightarrow a_{1}|\mu_{1}^{+}\rangle+a_{2}|\mu_{2}^{+}\rangle-d_{1}|\nu_{1}^{-}\rangle-d_{2}|\nu_{2}^{-}\rangle$ $=-b_{1}|\mu_{1}^{-}\rangle-b_{2}|\mu_{2}^{-}\rangle+c_{1}|\nu_{1}^{+}\rangle+c_{2}|\nu_{2}^{+}\rangle$ (S9)
$\Rightarrow\;\;\;\,\Bigg[|\mu_{1}^{+}\rangle,|\mu_{2}^{+}\rangle,-|\nu_{1}^{-}\rangle,-|\nu_{2}^{-}\rangle\Bigg]\begin{bmatrix}a_{1}\\ a_{2}\\ d_{1}\\ d_{2}\end{bmatrix}$ $=\Bigg[-|\mu_{1}^{-}\rangle,-|\mu_{2}^{-}\rangle,|\nu_{1}^{+}\rangle,|\nu_{2}^{+}\rangle\Bigg]\begin{bmatrix}b_{1}\\ b_{2}\\ c_{1}\\ c_{2}\end{bmatrix}$. (S10)
On the left, we group together the amplitudes for the modes growing/propagating toward the domain wall, and on the right we have the modes growing/propagating away from the domain wall. Inverting the matrix on the right, we arrive at a scattering matrix relation of the form
$S(k_{x},E,m_{0},\Delta\theta)\begin{bmatrix}a_{1}\\ a_{2}\\ d_{1}\\ d_{2}\\ \end{bmatrix}=\begin{bmatrix}b_{1}\\ b_{2}\\ c_{1}\\ c_{2}\\ \end{bmatrix}$. (S11)
Now we can vary $E$ and/or other parameters, and monitor the eigenvalues of $S$. A zero eigenvalue corresponds to the existence of an eigenstate of $H$ that is exponentially localized to $y=0$. (Note that propagating waves cannot yield zero eigenvalues due to energy conservation). In Fig. 1b of the main text, we plot the minimum squared eigenvalue, $\mathrm{min}[\mathrm{eig}(S^{\dagger}S)]$, versus $E$ and $k_{x}$ and/or $\Delta\theta$, and use the zeros of this quantity to locate the Jackiw-Rebbi-like bound states, which agree well with the previous analytical results.
S3 Numerical solutions of bound states at 2D phase dislocation
To solve the Jackiw–Rossi equations using the finite element method (FEM), we express the equations in weak form. We start with the time-independent Schrödinger equation
$\begin{bmatrix}0&-i\partial_{x}-\partial_{y}&m_{1}-im_{2}&0\\ -i\partial_{x}+\partial_{y}&0&0&m_{1}-im_{2}\\ m_{1}+im_{2}&0&0&i\partial_{x}+\partial_{y}\\ 0&m_{1}+im_{2}&i\partial_{x}-\partial_{y}&0\\ \end{bmatrix}\begin{bmatrix}\psi_{1}\\ \psi_{2}\\ \psi_{3}\\ \psi_{4}\end{bmatrix}=E\begin{bmatrix}\psi_{1}\\ \psi_{2}\\ \psi_{3}\\ \psi_{4}\end{bmatrix}$, (S12)
where $\psi_{1,2,3,4}$ are the wavefunction components and $E$ is the energy eigenvalue. Next, we multiply each component of this equation with a set of test functions $\phi_{1,2,3,4}$, and integrate over some area $\Omega$:
$\int_{\Omega}\phi_{1}\cdot\big[-i(\partial_{x}\psi_{2}-i\partial_{y}\psi_{2})+(m_{1}-im_{2})\psi_{3}\big]\;d^{2}r$ $=E\int_{\Omega}\phi_{1}\cdot\psi_{1}\;d^{2}r$ (S13)
$\int_{\Omega}\phi_{2}\cdot\big[-i(\partial_{x}\psi_{1}+i\partial_{y}\psi_{1})+(m_{1}-im_{2})\psi_{4}\big]\;d^{2}r$ $=E\int_{\Omega}\phi_{2}\cdot\psi_{2}\;d^{2}r$ (S14)
$\int_{\Omega}\phi_{3}\cdot\big[(m_{1}+im_{2})\psi_{1}+i(\partial_{x}\psi_{4}-i\partial_{y}\psi_{4})\big]\;d^{2}r$ $=E\int_{\Omega}\phi_{3}\cdot\psi_{3}\;d^{2}r$ (S15)
$\int_{\Omega}\phi_{4}\cdot\big[(m_{1}+im_{2})\psi_{2}+i(\partial_{x}\psi_{3}+i\partial_{y}\psi_{3})\big]\;d^{2}r$ $=E\int_{\Omega}\phi_{4}\cdot\psi_{4}\;d^{2}r.$ (S16)
We then apply integration by parts to the derivative terms on the left side of each equation. For example, for the first term in the left hand side of (S13), we have
$\int_{\Omega}\phi_{1}\cdot(-i\partial_{x}\psi_{2})\,d^{2}r=i\int_{\Omega}(\partial_{x}\phi_{1})\cdot\psi_{2}\,d^{2}r-i\oint_{\partial\Omega}\left(\phi_{1}\cdot\psi_{2}\right)n_{x}\;d\ell,$ (S17)
where $\Omega$ and $\partial\Omega$ are the domain and its boundary, $n_{x}$ is the $x$-component of outward normal vector on $\partial\Omega$, and $d\ell$ is the boundary element. The first term on the right of (S17) enters into our desired weak form of the PDE. The second term is a boundary term that is eliminated in the usual way (e.g., by imposing Dirichlet conditions on distant boundaries). Carrying out this procedure for all the terms in (S13)–(S16), we obtain
$\int_{\Omega}\big[+i(\partial_{x}\phi_{1})\psi_{2}+(\partial_{y}\phi_{1})\psi_{2}+(m_{1}-im_{2})\phi_{1}\psi_{3}-E\phi_{1}\psi_{1}\big]d^{2}r$ $=0$ (S18)
$\int_{\Omega}\big[+i(\partial_{x}\phi_{2})\psi_{1}-(\partial_{y}\phi_{2})\psi_{1}+(m_{1}-im_{2})\phi_{2}\psi_{4}-E\phi_{2}\psi_{2}\big]d^{2}r$ $=0$ (S19)
$\int_{\Omega}\big[-i(\partial_{x}\phi_{3})\psi_{4}-(\partial_{y}\phi_{3})\psi_{4}+(m_{1}+im_{2})\phi_{3}\psi_{1}-E\phi_{3}\psi_{3}\big]d^{2}r$ $=0$ (S20)
$\int_{\Omega}\big[-i(\partial_{x}\phi_{4})\psi_{3}+(\partial_{y}\phi_{4})\psi_{3}+(m_{1}+im_{2})\phi_{4}\psi_{2}-E\phi_{4}\psi_{4}\big]d^{2}r$ $=0$ (S21)
The integrands are the weak expressions entered into the FEM solver. The phase and amplitude of the complex mass are input as $m_{1}=|m(\textbf{{r}})|\cos(\theta(\textbf{r}))$ and $m_{2}=|m(\textbf{r})|\sin(\theta(\textbf{r}))$. The numerical solutions are obtained using Dirichlet boundary conditions, $\psi_{1,2,3,4}=0$, at the edges of the computational cell, far from the plotted region.
S4 Acoustic structure design
Figure S3: Design of the acoustic structure. a, Band structure for the acoustic crystal with Kekulé modulation (red lines, modulation phase $\theta=-0.35\pi$) and without it (gray lines). All other parameters are the same as in Fig. 2 of the main text. b Bulk band gap versus modulation phase $\theta$, with bulk bands shaded in gray. c, A phase map $\theta(\mathbf{r})$ with a phase singularity. d, Eigenmode spectrum for the acoustic structure generated from c. The state in the bulk gap is a Jackiw–Rossi or “Dirac vortex” state. Inset: spatial profile of the in-gap state, which is localized at the center of the phase singularity and is thus interpretable as the Jackiw–Rossi mode [21]. All results shown here are obtained from FEM simulations.
The acoustic structure that we use to realize the Jackiw–Rossi model, and the consequent DBC modes, is similar to the design in Ref. 44. We begin with a triangular lattice of solid pillars surrounded by air, as shown in Fig. 2a of the main text. If all pillars have equal radii, the band structure has two Dirac points at the corners of the hexagonal Brillouin zone. If the lattice is viewed in terms of an expanded unit cell with six pillars each, both Dirac points are folded to the center, $\Gamma$, of the shrunken Brillouin zone [8]. In Fig. S3a, the gray lines show the acoustic band structure calculated by FEM. Near $\Gamma$, therefore, we have four bands (two from each valley) governed by an effective Jackiw–Rossi model with $m_{1}=m_{2}=0$.
We then apply a Kekulé-type modulation by varying the pillar radii according to Eq. (6) of the main text, which involves a modulation amplitude $\Delta R\neq 0$ and phase $\theta$. The modulation lifts the Dirac point degeneracy, as seen in Fig. S3a (the red lines show the bands for $\theta=-0.35\pi$). Since the modulation phase parameter $\theta$ is an angle variable by construction, we infer that it maps to the phase parameter $\mathrm{arg}[m]$ for the complex mass field [44]. In fact, this parameter mapping is not exact: the modulation phase also induces a variation in $|m|$, as seen in the variation in the bulk gap in Fig. S3b. In the present work, we neglect the $|m|$ variation when designing our DBC waveguide structures; in the future, further optimizing the modulation scheme may enable creating more performant structures.
Thus, to realize a Jackiw–Rossi model with spatially-varying $\mathrm{arg}[m(\mathbf{r})]$, we apply the corresponding spatially-varying modulation phase parameter $\theta(\mathbf{r})$ [28, 30]. For instance, it is possible to generate a Jackiw–Rossi (or “Dirac vortex”) state with a $\theta(\mathbf{r})$ distribution like the one shown in Fig. S3c, containing a phase singularity. Using this $\theta(\mathbf{r})$, we generate the acoustic structure and calculate its eigenmodes; the results are shown in Fig. S3d, and indeed feature a localized eigenmode in the gap. In the main text, we proceed to study $\theta(\mathbf{r})$ distributions containing branch-cuts, which host DBC modes.
S5 Comparison to defect mode cavity
In the main text, we observed that the DBC modes can be used to form line cavities, by using a branch cut terminating at finite branch points. Similar line cavities can also be formed with conventional defect modes, by altering a line of unit cells in an otherwise uniform lattice [49]; typically, these altered cells are placed in straight line, since conventional defect modes are susceptible to Anderson localization in the longitudinal direction (i.e., along the cavity line). Aside from the Anderson localization issue, DBC modes may also possess an advantage in having more uniform transverse confinement lengths, as noted in the main text. Here, we use FEM simulations to perform such a comparison.
Figure S4: Comparison of DBC line cavity with conventional defect cavity. a, DBC mode spectrum in the acoustic structure, reproduced from Fig. 2c (left panel) of the main text. The red curves correspond to the DBC modes. b, Q-factors for the DBC line cavity eigenmodes. The line cavity design is the same as in Fig. 3 of the main text, but with length extended to 600 mm. c, Normalized transverse mode profiles for the four eigenmodes marked by stars in b. The gaps in the curves are due to overlaps of the measurement line with pillars. d, Spectrum of acoustic crystal defect modes in a square array of solid pillars. Inset: structure of the waveguide, which is formed by omitting one line of air holes, and slightly enlarging the radius of two side pillars by 1.2 times. The unmodulated pillar radius is r=r=7 mm, and the lattice constant is p=p=18.4 mm. e, Q-factors for the eigenmodes of a line defect cavity of length 360 mm. f, Normalized transverse mode profiles for the four eigenmodes marked by stars in e. In a–b and d–e, the bulk bands are shaded in gray.
For the DBC-based line cavity, we use the acoustic structure described in Sec. S4 and the main text, with $\Delta\theta=\pi$. The dispersion relation is shown in Fig. S4a. We use this to form a line cavity similar to Fig. 3a of the main text, except that the length is extended to 600 mm; all other design parameters are the same. Note this is a standalone line cavity, with no coupling waveguide. The simulated eigenmode Q-factors are plotted in Fig. S4b. We will focus on the four cavity eigenmodes around $10\,\textrm{kHz}$, which are formed by the lower band of DBC modes, because their Q-factors and proximity to the band gap edge are similar to the conventional cavity eigenmodes described below. In Fig. S4c, we plot transverse profile for these four eigenmodes, which reveals that their confinement lengths are all very similar—even the curves plotted in red and blue, which correspond to eigenmodes that lie very close to the band gap edge or just inside the bulk band (see Fig. S4b).
Next, we simulate a line cavity using conventional defect modes. The dispersion relation is shown in Fig. S4d; the acoustic structure consists of a square lattice of pillars with modified/omitted pillars, as shown in the inset of Fig. S4d. Upon forming a line cavity of length 360 mm, we obtain the eigenmodes shown in Fig. S4e. We focus on the four eigenmodes closest to the upper edge of the band gap, which have Q-factors similar to those in Fig. S4b. The corresponding transverse mode profiles are plotted in Fig. S4f, and exhibit noticeably larger variations in confinement length.
S6 Photonic analogue
In this section, we present FEM simulation results showing that the present scheme for realizing DBC modes in acoustics can be transferred to photonics.
In the photonic design, the acoustic structure’s pillars are converted into air holes, while the air regions are converted into GaAs. With a lattice constant of $a=28\mu\textrm{m}$ and GaAs refractive index of 3.74, the PhC supports a transverse magnetic (TM) band gap near 3.8 THz, as shown in Fig. S5a. For a straight DBC waveguide, analogous to the acoustic design of Fig. 2 in the main text, this yields the dispersion relation shown in Fig. S5c. Similar to the acoustic case, we can use the DBC modes to form various configurations such as line cavities, as shown in Fig. S5cd. Photonic realizations at higher frequency regimes can also be achieved by scaling down the dimensions of the structure.
Figure S5: Photonic DBC modes. a, Terahertz-scale photonic structure, similar to the acoustic design but with air holes in a GaAs medium. The lattice constant is aa=28 $\mu m$, and the refractive index of GaAs is 3.74. The domain wall is marked by dashes. b, Dispersion relations for the transverse magnetic (TM) photonic DBC modes, calculated from FEM simulations. c, Q-factors for the eigenmode of a DBC line cavity, constructed similarly to Fig. 3 of the main text and Fig. S4ac, with a cavity length 360 $\mu m$. d, Field amplitude ($|E_{z}|$) distributions for two of the eigenmodes in c.